# https://fudan-zvg.github.io/Efficient4D

### v2

@@ 4 Experiments
4.1 Implementation details
C4D [13]
Ours
GT
Figure 4: Qualitative evaluation of Consistent4D (C4D) [13] and our method against ground truth (GT) on synthetic data.
Competitors
Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [13]. In comparison, we obtained the results of Consistent4D by running its released code. We also compare SyncDreamer [22] partially by replacing SyncDreamer-T in Sec. 4.5.
Evaluation videos
@@ 4.1 Implementation details
To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. Specifically, we focused on 36 sequences: 32 sequences released by [13] and 4 sequences processed by ourselves. Among the released data, seven sequences are synthetic data where ground truth are provided. Our four sequences are used for sparse input evaluation in Sec. 4.4, which only contain two frames for each sequence. Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab [2]. The other one named yellow face was collected from the internet. Each sequence consists of monocular videos captured by a stationary camera, with duration ranging from approximately 1s to 3s.
Evaluation metrics
@@
As 4D generation research is still at early stage, there is no well established metric yet. However, we adopt multiple metrics by referring related works for comprehensive evaluation. For evaluating on synthetic data, we use LPIPS score [51] and CLIP similarity [32] between rendered images and ground truth following [13]. For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as [39, 33] to measure image quality. For temporal smoothness, following [50, 7, 9] we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views. To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
Implementation
For SyncDreamer-T, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$ in time-synchronous spatial volume. The improved sampling strategy [45] can be seamlessly integrated into SyncDreamer-T. In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian. Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 20,000. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 20,000 iterations with a batch size of 2. Gaussians are densified every 200 iterations, with densification halted at the 10,000th iteration. All other hyperparameters remain consistent with those in [15]. For speed efficiency, our proposed method only costs about 2 minutes for image generation and 4 minutes for reconstruction on one A6000 GPU.
4.2 Qualitative evaluation
Input samples
Refer to caption Refer to caption Refer to caption Refer to caption
Consistent4D [13] ~165 min
Refer to caption Refer to caption Refer to caption Refer to caption
Our images ~2 min
Refer to caption Refer to caption Refer to caption Refer to caption
Our geometry ~4 min
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparisons. We show four images per method, captured from two viewpoints at two timestamps. Our geometry is the points at 3D mean of Gaussians.
@@ 4.2 Qualitative evaluation
We first present the results on synthetic data in Fig. 4 with ground truth shown. Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth. For other data without ground truth, we present qualitative comparisons in Fig. 5. In geometry rendering, we extract points at the mean of Gaussians. We also eliminate white background points and transparent points with low opacity. When assessing texture quality, it is important to note that both methods fall under the category of lifting 2D to 4D. However, Consistent4D produces watercolor-like images with low fidelity, which can be attributed to conflicts from multiple supervisory signals during prolonged optimization. In contrast, our method excels in directly generating high-quality 2D images. Subsequently, our method’s reconstruction stage yields explicit point clouds and exhibits less degradation in rendering quality compared to the SDS-based Consistent4D. Overall, our method consistently outperforms Consistent4D in most cases.
4.3 Quantitative evaluation
Table 1: Quantitative comparisons. The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering. : Measured using code released by Consistent4D [13] with one A6000 GPU.
Consistent4D [13] Efficient4D(Ours)
- Comparison with ground truth on synthetic data
CLIP $\uparrow$ 0.87 0.92
LPIPS $\downarrow$ 0.16 0.14
- Comparison on all cases
CLIP $\uparrow$ 0.8471 0.8721
CLIP-T-f $\uparrow$ 0.9692 0.9702
CLIP-T-s $\uparrow$ 0.9658 0.9726
CLIP-T-b $\uparrow$ 0.9697 0.9707
Generation time $\downarrow$ 120 mins 6 mins
FPS $\uparrow$ ~20 ~1000
Explicit geometry
In Tab. 1, we compare Consistent4D and our method quantitatively. The metrics and data used are described in Sec. 4.1.
We draw several observations: (1) For synthetic data, our generation results produce better CLIP and LPIPS scores. (2) Our method achieves superior image quality and temporal smoothness on all cases. (3) Our method significantly accelerates the generation process, achieving over 20$\times$ speed improvement. (4) Our efficient 4D Gaussian representation and splatting based rendering enable real-time rendering and provide explicit point cloud geometry.
More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization faster, and (2) each iteration in our method requiring much less time compared to Consistent4D. This time efficiency is further enhanced by avoiding recurrent evaluation in the diffusion model and prolonged forward rendering of the inefficient 4D representation as needed in the competitor at every iteration.
@@ 4.3 Quantitative evaluation
4.4 Sparse input evaluation
We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames. In such cases, we set $w=0.25$ in Eq. 2. As illustrated in Fig. 6, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency. In contrast, Consistent4D fails to operate effectively under such conditions.
Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the video-to-4D task.
@@ 4.5 Ablation studies
K-Plane [8] PSNR: 33.50 Time: 2 hrs
Refer to caption Refer to caption
No conf PSNR: 32.13 Time: 6 mins
Refer to caption Refer to caption
Full setting PSNR: 32.64 Time: 6 mins
Refer to caption Refer to caption
Reconstructed novel views
Figure 8: Ablation study on 4D scene representation and confidence maps. The images follow a chronological order from left to right.
@@
We performed ablation studies to assess the influence of different components in our approach. We compared our full method against five baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded; (4) Confidence maps are excluded; (5) 4D Gaussian splatting is replaced with K-Plane [8]. Note setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer. The results corresponding to these settings can be found in Fig. 7 and Fig. 8. Table 2 also gives quantitative evaluations.
Importance of synthetic data.
@@
We compare our generated image matrix with utilizing only input video using the same proposed 4D Gaussian representation model. As shown in the first row of Fig. 7, when only relying on a single-view video, the model cannot produce any meaningful results for novel views. This indicates the importance of constructing proper training data.
Effect of time-synchronous spatial volume.
@@
We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and last rows in Fig. 7 and Tab. 2. Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry. In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation and higher CLIP-T scores.
Effect of frame interpolation.
@@
As illustrated in the contrast between the third and last rows in Fig. 7 and Tab. 2, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views, thus delivering higher CLIP-T scores. This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
Effect of confidence-aware in objective.
The integration of a confidence-aware loss in our design serves as a strategy to mitigate training data noise. This is evident in the contrast between the “No conf” and “Full setting” results in Fig. 8 and Tab. 2. The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality and temporal smoothness.
Integrating existing 4D representation model.
Our synthetic training data is versatile and supports the optimization of 4D representation models, such as K-Plane [8]. In the comparison shown in Fig. 8, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes. Notably, our method surpasses K-Plane in speed, achieving a 20$\times$ improvement. Additionally, as emphasized in Sec. 4.3, our model enables real-time rendering with explicit point cloud geometry.
@@ 5 Conclusion
This study introduces a new framework, Efficient4D, designed for generating efficiently dynamic 4D objects seamlessly from monocular videos captured by a stationary camera. The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions using an innovative 4D Gaussian splatting model. Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 20 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results. Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.
Limitations.
Our local smoothing approach in sliding window style with our image sequence generator poses challenges in handling long-duration videos. To address this issue, replacing the smoothing filtering layer with a learnable attention layer featuring global receptive fields is a viable solution, particularly when dealing with extensive 4D datasets. Additionally, handling long videos may necessitate significant GPU memory, which can be mitigated by employing multi-GPUs or CPUs, albeit with an increase in processing time.

### v3

@@ 5 Experiments
5.1 Experiment setup
[…]
Competitors For comparison, we mainly focus on video-to-4D and image-to-4D task. The competitors include Consistent4D [14] (ICLR2024), 4DGen [18] (ArXiv2024) and STAG4D [19] (ECCV2024) for video-to-4D, and Animate124 [50] (ArXiv2023) and DreamGaussian4D [51] (ArXiv2024) for image-to-4D. We obtained their results by running their released official code. We also compare SyncDreamer [8] partially by replacing SyncDreamer-T in ablation study (Section 5.6).
[…]
@@ 5.1 Experiment setup
Evaluation data To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. For video-to-4D, we focused on 36 sequences: 32 sequences released by [14] and 4 sequences processed by ourselves. Among the released data, seven sequences are synthetic data where ground truth are available. For image-to-4D, we used the ten images released by [50]. Our four sequences are used for sparse input evaluation in Section 5.5, which only contain two frames for each sequence. Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab [52]. The other one named yellow face was collected from the internet. All the data are publicly available.
[…]
@@
Evaluation metrics As 4D generation research is still at early stage, there is no well established metric yet. However, we adopt multiple metrics by referring related works for comprehensive evaluation. For evaluating on synthetic data, we use LPIPS score [53] and CLIP similarity [54] between rendered images and ground truth following [14]. For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as [55, 51] to measure image quality. For temporal smoothness, following [18, 56, 57] we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views. To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
[…]
Implementation details In equation (8), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase. In practice, we can adjust the weight following two rules: (i) the middle weight should be at least $0.5$; (ii) At least half of the volumes should have a positive weight. Here we provide some sample weights which have a good effect:
$\mathbf{w}=\left\{\begin{matrix}(1.5,7,1.5),\quad 3\leq N<5\\ (1,1,6,1,1),\quad 5\leq N<7\\ (1,1,1,6,1,1,1),\quad 7\leq N<13\\ (1,1,1,1,1,2,14,2,1,1,1,1,1),\quad N\geq 13,\end{matrix}\right.$ (14)
where $N$ denotes the frame number. The weight will be normalized to ensure the sum is $1$.
In denosing process, we follow the default setting of [8] using a improved sampling strategy introduced by HarmonyView [58]. HarmonyView redefines the score estimation by decomposing consistency and diversity. Please refer to [58] for details.
In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 5,000. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 1500 iterations with batch size 1. All other 4Dgs hyperparameters remain consistent with those in [15]. The balancing weights in loss function (equation 13) are set as $\tilde{\lambda}_{\text{rgb}}=8000,\tilde{\lambda}_{\text{ssim}}=2000,\lambda_{\text{SDS}}=1$. In equation (12), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model [48]. For each generated frame $I_{t}$, we use its adjacent four frames to interpolate the estimated frame as follows:
$\hat{I}_{t1}=\text{Interp}(I_{t-1},I_{t+1}),~{}~{}\hat{I}_{t2}=\text{Interp}(I_{t-2},I_{t+2}).$ (15)
Then we can compute the mean RGB confidence score:
$\mathcal{C}_{\text{rgb}}=1-\frac{1}{2}\left(|I_{t}-\hat{I}_{t1}|+|I_{t}-\hat{I}_{t2}|\right).$ (16)
The SSIM confidence score is similar by using equation (12).
For speed efficiency, our proposed method only costs about 2 minutes for image generation if parallel denoising is allowed and 8 minutes for reconstruction on one A6000 GPU.
5.2 Evaluation on synthetic data
Figure 4: Qualitative evaluation against ground truth (GT) on synthetic data. We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19].
Table 1: Quantitative evaluation on synthetic data. We report CLIP and LPIPS scores between rendered images and ground truth images.
Consistent4D [14] 4DGen [18] STAG4D [19] Efficient4D(Ours)
CLIP $\uparrow$ 0.87 0.89 0.91 0.92
LPIPS $\downarrow$ 0.16 0.14 0.13 0.13
We first present the results on synthetic data in Figure 4 with ground truth shown. Since the data only include videos, we compare with two video-to-4D methods: Consistent4D [14], 4DGen [18] and STAG4D [19]. Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth, such as the direction of dinosaur’s head, the color of trump’s arm and the teeth of skull. The quantitative metrics are also presented in Table 1. The better CLIP and LPIPS scores further validate our superior generation results.
5.3 Video-to-4D comparison
Figure 5: Qualitative comparisons on video-to-4D generation. We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19]. For each case, we show four images per method with $0^{\circ}$ elevation. Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
Figure 6: Qualitative comparisons on image-to-4D generation. We compare our Efficient4D with Animate124 [50] and DreamGaussian4D [51] (DG4D). For each case, we show four images per method with $0^{\circ}$ elevation. Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
Table 2: Quantitative comparisons with state-of-the-art methods on both video-to-4D and image-to-4D generation.
[…]
- Video-to-4D comparison
Consistent4D [14] 0.8471 0.9692 0.9658 0.9697 120 mins
4DGen [18] 0.8730 0.9568 0.9568 0.9573 130 mins
STAG4D [19] 0.8398 0.9766 0.9731 0.9760 70 mins
Efficient4D (Ours) 0.8745 0.9766 0.9609 0.9780 10 mins
- Image-to-4D comparison
Animate124 [50] 0.8076 0.9673 0.9639 0.9541 540 mins
DreamGaussian4D [51] 0.8242 0.8999 0.9072 0.9038 12 mins
Efficient4D (Ours) 0.8350 0.9346 0.9297 0.9321 10 mins
@@ 5.3 Video-to-4D comparison
For data without ground truth, we present qualitative comparisons in Figure 5. When assessing texture quality, it is important to note that all methods fall under the category of lifting 2D to 4D. However, Consistent4D produce watercolor-like images with low fidelity, which can be exemplified by the blurry edges of their rendered images. We attribute this blur to conflicts from multiple supervisory signals during prolonged optimization. In contrast, our method excels in directly generating high-quality 2D images, thus providing a strong and consistent supervision in the reconstruction stage. Although 4DGen and STAG4D also use pseudo labels, their images are inconsistent in the temporal coordinate and SDS loss still dominates the optimization. Therefore, 4DGen lacks details and STAG4D suffers from floaters and blur. Besides, our method’s reconstruction stage supports rendering under any viewpoints while the images generated in stage-1 are discrete and sparse with a fixed elevation $30^{\circ}$. Overall, our method consistently outperforms the baseline methods in most cases.
In Table 2, we compare our method with baselines quantitatively. The metrics and data used are described in Section 5.1. We draw several observations: (i) Our method achieves superior image quality and temporal smoothness on all cases validated by the higher CLIP and CLIP-T scores; (ii) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement (10 mins vs 120 mins). More specifically, the speed improvement is attributed to (i) our image supervision design, converging optimization faster and reduce the training iterations, and (ii) each iteration in our method requiring much less time thanks to the efficiency of Gaussian representation.
5.4 Image-to-4D comparison
Next, we will evaluate our method on the image-to-4D task. In Figure 6, we present the visual results of three different methods. It can be observed that our extended method is capable of generating 4D assets with higher quality compared to those produced by state-of-the-art image-to-4D methods. Animate124 generates textures lacking in detail, and DreamGaussian4D tends to produce fragmented meshes. In contrast, the results of our Efficient4D exhibit both good geometry and high-quality textures. Additionally, Table 2 supports the similar conclusion that our method is superior, as seen in the video-to-4D comparison. Although Animate124 achieves the best CLIP-T score, This score merely indicates that the range of motion is small rather than the motion being continuous. From Figure 6 we can easily find that the quality of Animate124 results are lower than any other methods. Another notable point is that the overall CLIP scores for image-to-4D methods are lower than those for video-to-4D methods. This may be because the reference single-view videos in image-to-4D are generated by video diffusion models, which are inconsistent and have lower quality compared to real-world videos.
5.5 Sparse input case
@@
We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames. In such cases, we set $w=0.25$ in equation (9). We also modified the code of Consistent4D [14] to make it adapt to such case. As illustrated in Figure 7, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency. In contrast, Consistent4D fails to operate effectively under such conditions.
Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
@@ 5.6.1 Ablation of image generation
We first performed ablation studies to assess the influence of different components in our image generation stage. To better evaluate the image quality itself, we reconstruct the 4D Gaussian without SDS loss. We compared our full method against three baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded. Note that setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer. The results corresponding to these settings can be found in Figure 8. Table 3 also gives quantitative evaluations.
[…]
@@
Importance of synthetic data We compare our generated image matrix with utilizing only input video using the same 4D Gaussian representation model. As shown in the first row of Figure 8, when only relying on a single-view video, the model cannot produce any meaningful results for novel views. This indicates the importance of constructing proper training data.
[…]
@@
Effect of time-synchronous spatial volume We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and last rows in Figure 8. Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry. In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation and higher CLIP-T scores in Table 3.
[…]
@@
Effect of frame interpolation As illustrated in the contrast between the third and last rows in Figure 8, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views, thus also delivering higher CLIP-T scores in Table 3. This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
5.6.2 Ablation of reconstruction stage
Figure 9: Ablation study on confidence maps. The images follow a chronological order from left to right.
Figure 10: Ablation study on image supervision and SDS loss.
In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, and (3) SDS loss. The visual results are shown in Figure 9 and 10.
[…]
Effect of confidence map The integration of a confidence-weighted loss in our design serves as a strategy to mitigate training data noise. In the left of Figure 9, we study the effect of confidence map when obvious inconsistency exists in the generated images. Our temporal smooth mechanism can support consistency for most areas, whilst a small fraction of lower-quality regions may introduce conflicting gradients in the reconstruction stage, thus hurting the overall quality. To mitigate this, our confidence aware design comes into play for weakening the supervision from those inconsistent regions. Since these weakly supervised regions only occupy a small proportion across all views, the missing information about texture or geometry can be compensated by the redundancy of other views and generalization ability of 4D Gaussians. As shown in the left of Figure 9, the inclusion of confidence maps effectively reduces blurry and inconsistent rendering even when the generated images have plausible consistency, resulting in a significant overall improvement in quality and temporal smoothness.
[…]
Importance of image supervision and SDS loss In Figure 10, we compare the full setting of our method with baselines without image supervision (w/o image) or SDS loss (w/o SDS). The absence of anchored image supervision results in bad geometry, such as the hole in the elephant ears and multiple legs of the lion. This can be also attributed to the conflicts of SDS loss during prolonged optimization. In addition, due to the sparsity of the generated images, the novel views may not be reconstructed well, leading to blur and floaters. By contrast, our full method avoids the disadvantages of both, being able to render clean images with good geometry.
@@ 6 Conclusion
This study introduces a new framework, Efficient4D, designed for generating efficiently dynamic 4D objects seamlessly from monocular videos captured by a stationary camera. The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions. Our approach, utilizing image supervision with lightly weighted SDS loss, significantly accelerates the generation process, achieving about 10 times faster speeds compared to previous works, while still delivering superior reconstruction and novel view synthesis results. Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.